OpenStudy (anonymous):

find the vector T at the given point r(t) = ; point (1,2/3,1) I do not know how to count t at the point. Please, help

OpenStudy (phi):

I don't understand the question. As far as I can see if (1, 2/3, 1) is on r(t), then we can equate corresponding values 1 = t^2 2/3 = 2/3 t^3 1 = t in each case, t=1 (which is good, otherwise it doesn't make sense) but what is vector T ?

OpenStudy (phi):

is vector T the tangent ?

OpenStudy (phi):

if so, take the derivative wrt to t, and sub in 1 for t in the derivative.